40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Tue Feb 9 22:43:29 2021, rana, Noise Hunting, Laser, "North" current driver having slow current fluctuations PXL_20210210_063043141.jpg
    Reply  Wed Feb 10 09:52:45 2021, shruti, Noise Hunting, Laser, "North" current driver having slow current fluctuations 
    Reply  Mon Feb 15 21:25:00 2021, rana, Noise Hunting, Laser, "North" current driver having slow current fluctuations Planex-Beatnote.pdfPlanex-Beatnote-ASD.pdfplotPhaseMeter.tar.gz
    Reply  Thu Feb 18 12:30:09 2021, aaron, Noise Hunting, Laser, "North" current driver having slow current fluctuations 
Message ID: 2651     Entry time: Thu Feb 18 12:30:09 2021     In reply to: 2636
Author: aaron 
Type: Noise Hunting 
Category: Laser 
Subject: "North" current driver having slow current fluctuations 

I dug through the log to figure out what created the 8s noise in the current driver. I suspect it was my power cable connectors on Feb 3 , but it seems Rana measured this clean beatnote spectrum on Feb 4. Still, all other changes were downstream of the current drives -- was the broken current driver involved in that clean beat? 

I've added another datapoint of the laser frequency noise from my measurement Jan 25.

ELOG V3.1.3-